# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v1

@@ Abstract
As artificial intelligence becomes increasingly sophisticated, companion chatbots have been proposed as a potential solution to the growing epidemic of loneliness. However, the impact of these AI companions on users’ psychological well-being and social behaviors remains poorly understood. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and their experience of loneliness. We find a small but significant direct correlation between session length with chatbots and loneliness, and we develop a model of this relationship through multiple regression analysis, finding social attraction and neuroticism as moderators. We find seven clusters of users, including socially fulfilled dependent users to lonely dependent users. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for what kind of human-AI connections might lead to emotional well-being and complement rather than replace human connections.
@@ 1 Introduction
In an increasingly digital world, we experience both increased connectivity and increased social isolation. Despite unprecedented technological means to connect, loneliness remains a pervasive issue, affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness” in May 2023 [Office of the Assistant Secretary for Health (OASH)(2023)]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Chawla et al.(2021)].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, chatbots engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [Patel(2024), Shewale(2024)]. These AI companions offer users the opportunity for constant, judgment-free interaction. However, the impact of these artificial relationships on users’ psychological well-being and social behaviors needs further study.
Some research suggests that AI companions, even when used over the short term, decreases feelings of loneliness [Loveys et al.(2019), Gasteiger et al.(2021), De Freitas et al.(2024)]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [Turkle(2011)]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about the factors that determine whether and to what extent companion chatbots impact loneliness and other psychological and social factors. As Xia et al. [Xia et al.(2024)] noted, artificial agents may complement human social interactions in some contexts while replacing them in others. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being.
To address this gap, we conducted a large-scale survey (n = 404) of regular companion chatbot users, with the following research questions:
@@
Our study (see Figure 1) aims to provide a more nuanced understanding of the impact of companion chatbots, moving beyond binary assessments of positive or negative effects. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (such as reduced loneliness, improved social skills, and better mental well-being) and those who may experience negative consequences (such as emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of AI companions influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
The ethical implications of widespread AI companion use are multifaceted. While these technologies may serve as a valuable intervention for individuals with limited access to social support, there are concerns about the long-term effects on social skills and expectations for human relationships. Our study aims to provide empirical grounding for these discussions, informing both design practices and policy considerations in the rapidly evolving landscape of AI-mediated social interaction.
@@
  • A model for understanding the relationship between companion chatbot usage and loneliness, accounting for mediators such as socialization and moderators of social attraction and neuroticism.
  • A typology of companion chatbot users that takes into account usage patterns, psychological characteristics, and perceived benefits or drawbacks.
  • An exploration of motivations and usage characteristics of users of companion chatbots.
  • Implications for creating more effective and ethically responsible AI companions that enhance well-being.
@@ 2.1 Loneliness and Social Support
Loneliness, defined as the subjective feeling of isolation regardless of one’s actual social network, has become a pressing global issue affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. The U.S. Surgeon General’s declaration of loneliness as an epidemic in 2023 underscores the gravity of this issue [Office of the Assistant Secretary for Health (OASH)(2023)]. Chronic loneliness is associated with a significantly increased risk of mortality—approximately 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Freedman and Nicolle(2020)].
Various risk factors contribute to social isolation and loneliness, including socio-demographic factors (e.g., increased age, low income), medical factors (e.g., cognitive impairment, sensory loss), and social factors (e.g., living alone, lack of transportation) [Freedman and Nicolle(2020)]. Older populations are particularly vulnerable due to the accumulation of these risk factors and potentially lower digital literacy, which can limit their use of modern communication technologies [Gasteiger et al.(2021)].
@@ 2.3 Companion Chatbots
The emergence of companion chatbots, also known as virtual friends or digital companions, represents a novel approach to tackling loneliness [Wang et al.(2024), Goodings et al.(2024), Boine(2023), Chen et al.(2024), Shani et al.(2022), Brewer(2022)]. We refer to these as commercially available chatbots that employ generative AI to engage in complex conversations and that maintain persistent relationships with users. These companion chatbots are accessible to a broad demographic, with several popular companion chatbot services (see Figure 2) having gained significant user bases:
@@ 2.4 Human-AI Interaction for Emotional Support
Certain populations, such as LGBTQ individuals who may lack support from family and friends, are increasingly turning to LLM-based chatbots for mental health support. These users often seek out chatbots for immediately accessible support, viewing conversations with AI as safe spaces for emotional expression, showing the potential for AI to provide a sense of connection and safety for marginalized groups [Ma et al.(2024)].
@@
Neves et al. [Neves et al.(2015)] examined the difficulties faced by older adults in adopting new communication technologies, highlighting the importance of user-centered design and support systems in facilitating technology acceptance. These insights are crucial when considering the potential of AI chatbots as a scalable solution for emotional support and loneliness reduction.
While companion chatbots offer promising avenues for emotional support, there are significant ethical considerations and potential risks associated with their use. These include the possibility of over-reliance on AI, consequences for mood and affect, delays in seeking professional help, and further withdrawal from human socialization [Pataranutaporn et al.(2021)]. The development of dependence on AI agents and the potential replacement of real human relationships raise important questions about the long-term societal impacts of widespread AI companion use.
The tensions between the benefits and risks of AI companionship underscore the need for a nuanced understanding of how different users interact with and are affected by these technologies. As AI continues to advance, it becomes increasingly important to investigate the factors that determine positive or negative outcomes from companion chatbot use. This aligns with the principles of positive computing in human-AI interaction, which aims to design technology that supports psychological well-being [Calvo and Peters(2017)]. Towards this goal, our research seeks to inform the design of more effective and ethically responsible AI companions that support rather than supplant human social connections.
More specific, granular information about users and their needs is needed. Our research seeks to identify the characteristics of users who benefit from chatbot interactions, as well as those who may experience negative consequences. Our intent is to inform more targeted and ethical design practices in the rapidly evolving landscape of human-AI interaction.
@@ 3 Methods
We surveyed companion chatbot users recruited from CloudResearch, administering several psychological scales and measures of chatbot usage frequency and session length. Using the Qualtrics sample size estimator, we roughly estimated a minimum size of 385 with a 95% confidence level and a 5% margin of error. Participants were first recruited with a screening survey in which they were asked about whether or not they use companion chatbots regularly (at least once a week for at least a month). They were asked to elaborate on their experiences with chatbots briefly. Participants could only access the screening survey on CloudResearch if they were fluent in English and had prior experience using AI. A researcher manually read through the respondents’ brief free responses to assess whether they were engaging with companion chatbots regularly for social and emotional purposes. Those who were determined to be regular users were approved and given access to the full survey. Participants were screened and recruited through multiple rounds until we reached our target pool of qualified individuals. All qualifying participants were then invited to complete the main survey. We intentionally over-recruited to account for incomplete responses. After filtering out incomplete responses from the main survey, the final sample consisted of 404 valid responses (n = 404), slightly exceeding our initial target of 400. This study was approved by the institution’s IRB.
@@
  • Chatbot Usage Characteristics: We calculated descriptive statistics on multiple-choice questions regarding companion chatbot usage and examined participants’ free responses, both overall and grouped by the clusters we found. Our intention was to characterize our participants overall and provide more specific characteristics and examples of responses for each cluster.
@@
Figure 3: Comprehensive flowchart detailing the multi-method approach employed in this study. The diagram outlines the sequence of analytical steps, including participant recruitment, survey administration, data processing, correlation analysis, multiple regression modeling, cluster analysis, and the interpretation of both quantitative results and free-text responses.
@@ 3.1.1 Established Scales and Adaptations
We included several established scales, many of which were adapted to better fit our study context and to manage survey length:
  • UCLA Loneliness Scale (ULS-8) [Hays and DiMatteo(1987)]: Measures loneliness.
  • Lubben Social Network Scale (LSNS-6) [Lubben et al.(2006)]: Measures social isolation.
  • Multidimensional Scale of Perceived Social Support (MSPSS) [Zimet et al.(1990)]: Assesses perceived social support.
  • Brief Rosenberg Self-Esteem Scale (B-RSES) [Monteiro et al.(2022)]: Evaluates self-esteem.
@@
  • Big Five Inventory (BFI-10) [Rammstedt and John(2007)]: Evaluates personality traits based on five factors: extraversion, neuroticism (tendency to towards anxiety and other negative feelings), agreeableness (tendency to be compassionate and cooperative), conscientiousness (tendency to be self-disciplined and goal-directed), and openness to experience (tendency to be open-minded, imaginative, and insightful) [McCrae and Costa(1999)].
@@
  • Attitudes Towards AI Scale [Schepman and Rodway(2020)]: Measures attitudes towards artificial intelligence. Used the adaptation found in [Pataranutaporn et al.(2023)].
@@
  • Human-Chatbot Interaction Effect Scale [Xia et al.(2024)]: Measure of the effect of human-chatbot interaction on human-human interaction. Measures the respondent’s extraversion, expectations for chatbot services (in performance and meeting their needs), trust in interactions with social chatbots, perceived value of substituting people with chatbots, trust in interactions with people, and interpersonal communication intention (motivation and willingness to interact with people). Changed phrasing to refer to only chatbots and people rather than service providers. Removed the extraversion subscale, shortened subscales based on factor loadings, and made two versions of the interpersonal communication intention subscales, one for familiar people and one for unfamiliar people.
  • Generalized Problematic Internet Use Scale (GPIUS-2) [Caplan(2010)]: Shortened to 6 items based on relevance and subscale representation. Rephrased to reference chatbot use instead of general internet use.
  • State Empathy Scale [Shen(2010)]: Adjusted wording to be more general for application to both chatbots and people.
@@
  • Interpersonal Attraction Scale [McCroskey and McCain(1974)]: A scale measuring the respondent’s social attraction (desire to have a friendship and spend time with another), physical attraction (attraction to appearance), and task attraction (inclination to work with another based on competence and dependability). Shortened each of the three subscales to four items based on factor loadings and made two versions, one for chatbots and one for a close person; the version for a close person removed the physical attraction subscale due to the potential of the respondent’s close person being a family member.
@@
  • Perceived Homophily Scale [Mccroskey et al.(1975)]: A measurement of perceived similarity in interpersonal communication, which is associated with increased communication effectiveness. Changed to unidirectional Likert statements (e.g., “like me” instead of “unlike me – like me”). Selected 6 items from potential items based on factor loadings. Rephrased slightly into phrases and adjusted phrasing to make two versions, one for chatbots and one for a close person.
@@
  • Attributional Confidence Scale [Gudykunst and Nishida(1986)]: A measure of the respondent’s ability to predict others’ attitudes, feelings, and behavior [Clatterbuck(1979)]. Shortened to 5 items, removing items irrelevant to chatbots.
  • System Usability Scale (SUS) [Brooke(1996)]: A measurement of the respondent’s subjective assessment of a system’s usability. Adjusted phrasing to refer specifically to chatbots.
These scales were selected to provide a comprehensive view of participants’ social and psychological characteristics that might mediate the relationship between chatbot usage and loneliness. The adaptations were made to ensure relevance to the chatbot context and to manage survey length, aiming to reduce participant fatigue given the comprehensive nature of our investigation.

### v2

@@ Abstract
Companion chatbots offer a potential solution to the growing epidemic of loneliness, but their impact on users’ psychosocial well-being remains poorly understood. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and loneliness. We develop a model explaining approximately 50% of variance in loneliness; while usage does not directly predict loneliness, we identify factors including neuroticism, social network size, and problematic use. We identify seven distinct clusters of users, from socially fulfilled dependent users to lonely moderate users. Different usage patterns can lead to markedly different outcomes, with some users experiencing enhanced social confidence while others risk further isolation. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for developing more targeted and ethical approaches to AI companionship that complement rather than replace human connections.
@@ 1 Introduction
In an increasingly digital world, we are paradoxically more connected and isolated than ever. Despite unprecedented technological means to connect, loneliness remains pervasive, affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness”[Office of the Assistant Secretary for Health (OASH)(2023)]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day [Chawla et al.(2021)].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, which engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [Patel(2024), Shewale(2024)]. These AI companions offer users the opportunity for constant, judgment-free interaction. However, the impact of these artificial relationships on users’ psychological well-being and social behaviors needs further study.
Some research suggests that AI companions, even when used over the short term, decrease feelings of loneliness [Loveys et al.(2019), Gasteiger et al.(2021), De Freitas et al.(2024)]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [Turkle(2011)]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about factors influencing companion chatbots’ impact on loneliness and other psychological and social factors. As Xia et al. [Xia et al.(2024)] noted, artificial agents may complement human social interactions in some contexts while replacing them in others. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being.
To address this gap, we conducted a survey (n = 404) of regular companion chatbot users, with the following research questions:
@@
Our study (see Figure 1) aims to provide a more nuanced understanding of the impact of companion chatbots. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (i.e. reduced loneliness and improved social skills) and those who may experience negative consequences (i.e. emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of AI companions influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
@@
  • A model explaining approximately 50% of variance in loneliness, identifying problematic use as a key mediator and social attraction as a moderator in the relationship between chatbot usage and loneliness while accounting for personality traits and social network characteristics.
  • A typology of seven distinct companion chatbot user profiles demonstrating how similar usage patterns can lead to markedly different outcomes based on individual characteristics and social contexts.
  • An exploration of motivations and usage characteristics across different user types, revealing how technological exploration and entertainment dominate over social purposes despite marketing as social tools.
  • Design implications for creating more effective and ethically responsible AI companions, including the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns.
@@ 2.1 Loneliness and Social Support
Loneliness, defined as the subjective feeling of isolation regardless of one’s actual social network, has become a pressing global issue affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. Social isolation is the corresponding objective measure, often defined as a poorness in quantity and quality of contact with others. The U.S. Surgeon General’s declaration of loneliness as an epidemic in 2023 underscores the gravity of this issue [Office of the Assistant Secretary for Health (OASH)(2023)]. Chronic loneliness is associated with a significantly increased risk of mortality—approximately 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Freedman and Nicolle(2020)].
Various risk factors contribute to social isolation and loneliness, including socio-demographic factors (e.g., greater age, low income), medical factors (e.g., cognitive impairment, sensory loss), and social factors (e.g., living alone, lack of transportation) [Freedman and Nicolle(2020)]. While there is evidence of loneliness peaking in adolescence in certain demographics [Yang and Victor(2011)], older populations are particularly vulnerable due to the accumulation of these risk factors and potentially lower digital literacy, which can limit their use of modern communication technologies [Gasteiger et al.(2021)].
@@ 2.3 Companion Chatbots
The emergence of companion chatbots, also known as virtual friends or digital companions, represents a novel approach to tackling loneliness [Wang et al.(2024), Goodings et al.(2024), Boine(2023), Chen et al.(2024), Shani et al.(2022), Brewer(2022)]. We refer to these as commercially available chatbots that employ generative AI to engage in complex conversations and maintain persistent relationships with users. These companion chatbots are accessible to a broad demographic, with several popular companion chatbot services (see Figure 2) having gained significant user bases:
@@ 2.4 Human-AI Interaction for Emotional Support
Skjuve et al. described how human-chatbot relationships (HCRs) evolve from an initial superficial phase through more affective exploration – with greater trust and self-disclosure – to a more stable state of emotional connection [Skjuve et al.(2021)]. This relationship development enables companion chatbots to provide various forms of social support, including companionship, emotional support, and informational guidance in a “safe space” that is always available but not physical [Ta et al.(2020)]. The availability and perceived safety of these interactions make them an attractive alternative for social support for those who may lack support from family and friends, such as LGBTQ individuals [Ma et al.(2024)]. The motivations behind these interactions are diverse, ranging from curiosity to explicit seeking of emotional support, with users engaging in topics from casual conversation to deep personal issues. These findings suggest that while chatbots can serve as valuable sources of emotional support, the nature and impact of these interactions vary significantly based on individual user characteristics and needs [Ta-Johnson et al.(2022)]. Though limited to one service, Siemon et al.’s analysis of Replika reviews [Siemon et al.(2022)] showed that users of Replika often use the service to cope with their loneliness, and their long-term interactions with Replika can be joyful and beneficial.
@@
Age can also affect how users interact with AI. Neves et al. [Neves et al.(2015)] examined the difficulties faced by older adults in adopting new communication technologies, highlighting the importance of user-centered design and support systems in facilitating technology acceptance. These insights are crucial when considering the potential of AI chatbots as a scalable solution for emotional support and loneliness reduction.
While companion chatbots offer promising avenues for emotional support, there are significant ethical considerations and potential risks associated with their use. These include the possibility of over-reliance on AI, consequences for mood and affect, delays in seeking professional help, and further withdrawal from human socialization [Pataranutaporn et al.(2021)]. A recent case of a teen suicide after extensive use of Character.ai raised concerns regarding whether or not the chatbot could be blamed for his death [Roose(2024)]; this case underscores the urgent need to understand the influence of companion chatbots on psychosocial well-being. The development of dependence on AI agents and the potential replacement of real human relationships raise important questions about the long-term societal impacts of widespread AI companion use.
More specific, granular information about users and their needs is needed. Our research seeks to identify the characteristics of users who benefit from chatbot interactions and users who may experience negative consequences. Our intent is to inform more targeted and ethical design practices in the rapidly evolving landscape of human-AI interaction towards the goal of positive computing in human-AI interaction – designing technology that supports psychological well-being [Calvo and Peters(2017)].
@@ 3 Methods
We surveyed companion chatbot users recruited from CloudResearch, administering several psychological scales and measures of chatbot usage frequency and session length. This study was approved by the institution’s IRB. Using the Qualtrics sample size estimator, we roughly estimated a minimum size of 385 with a 95% confidence level and a 5% margin of error. Our criteria for participants were that they must be:
  • 18 years old or older
  • fluent in English
  • regular users of companion chatbots (using them at least once a week for at least a month).
Participants were first recruited with a screening survey, only accessible on CloudResearch if they were fluent in English and had prior experience using AI. A researcher manually read through respondents’ brief free responses regarding their experience with companion chatbots to remove participants who used chatbots only for practical purposes (i.e. assisting with coding or interview preparation) or those who did not describe any personal experiences (i.e. only describing the service). Participants were screened and recruited through multiple rounds until we reached our target pool of qualified individuals. All qualifying participants were then invited to complete the main survey. We intentionally over-recruited to account for incomplete responses. After filtering out incomplete responses from the main survey, the final sample consisted of 404 valid responses (n = 404), slightly exceeding our initial target of 400.
@@
  • Chatbot Usage Characteristics: We calculated descriptive statistics on multiple-choice questions regarding companion chatbot usage and examined participants’ free responses, both overall and grouped by the clusters we found. Our intention was to characterize our participants overall and provide more specific characteristics and examples of responses for each cluster. We also conducted thematic analysis of participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation using a stratified random sample of 15 participants per cluster (total N = 105).
@@
Figure 3: A flowchart detailing the multi-method approach employed in this study. The diagram outlines the sequence of analytical steps, including participant recruitment, survey administration, data processing, correlation analysis, multiple regression modeling, cluster analysis, and the interpretation of both quantitative results and free-text responses.
@@ 3.1.1 Established Scales and Adaptations
Our scale selection was guided by two principles: (1) measuring our primary variable of interest, loneliness, with a well-validated instrument, and (2) exploring potential mediating and moderating factors through an exploratory approach. While we acknowledge that including many scales risks multicollinearity, we addressed this through statistical controls in our analysis, detailed in the Analysis section.
Our primary dependent variable measure was the UCLA Loneliness Scale (ULS-8) [Hays and DiMatteo(1987)], demonstrating high internal consistency ($\alpha$ = .84) and construct validity through correlations with depression and inverse correlations with satisfaction with life, and social support [Wiseman et al.(1995), Wu and Yao(2008)]. It was selected for its brevity while maintaining psychometric properties of the original scale.
For our exploratory investigation of potential mediating and moderating factors, we included:
  • Lubben Social Network Scale (LSNS-6) [Lubben et al.(2006)]: Measures social isolation with strong internal consistency ($\alpha$ = .83). Selected to provide an objective measure of social connection to complement the subjective measure of loneliness.
  • Multidimensional Scale of Perceived Social Support (MSPSS) [Zimet et al.(1990)]: Assesses perceived social support. Selected for strong psychometric properties ($\alpha$ = .88) and wide usage [Wang et al.(2021)], as well as the known correlation between loneliness and social support [Ozbay et al.(2007), Wang et al.(2018), Schutter et al.(2021)].
  • Brief Rosenberg Self-Esteem Scale (B-RSES) [Monteiro et al.(2022)]: Evaluates self-esteem, a known risk factor leading to and increased by loneliness [Vanhalst et al.(2013)]. The brief scale was chosen for its shorter length but good internal consistency ($\alpha$ = .89 in a general population) and strong validity comparable to the original scale [Monteiro et al.(2022)].
@@
  • Big Five Inventory (BFI-10) [Rammstedt and John(2007)]: Evaluates personality traits based on five factors: extraversion, neuroticism (tendency to towards anxiety and other negative feelings), agreeableness (tendency to be compassionate and cooperative), conscientiousness (tendency to be self-disciplined and goal-directed), and openness to experience (tendency to be open-minded, imaginative, and insightful) [McCrae and Costa(1999)]. It correlates highly with the full BFI (mean correlation .83) [Rammstedt and John(2007)]. Selected based on previous research linking personality traits to loneliness [Buecker et al.(2020)].
@@
  • Attitudes Towards AI Scale [Schepman and Rodway(2020)]: Measures attitudes towards artificial intelligence with strong internal consistency for both the positive ($\alpha$ = .88) and negative ($\alpha$ = .83) subscales. Used the 8-item adaptation of the 16-item positive subscale, found in [Pataranutaporn et al.(2023)].
@@
  • Human-Chatbot Interaction Effect Scale [Xia et al.(2024)]: Measures the effect of human-chatbot interaction on human-human interaction. We included measures of the respondent’s expectations for chatbot services ($\alpha$ = .85), trust in interactions with social chatbots ($\alpha$ = .86), perceived value of substituting people with chatbots ($\alpha$ = .91), trust in interactions with people ($\alpha$ = .93), and motivation and willingness to interact with people ($\alpha$ = .94). Changed phrasing to refer to only chatbots and people rather than service providers, shortened subscales based on factor loadings, and made two versions of the interpersonal communication intention subscales, one for familiar people and one for unfamiliar people.
  • Generalized Problematic Internet Use Scale (GPIUS2) [Caplan(2010)]: A validated measure of problematic internet use with strong psychometric properties ($\alpha$ = .91 for full scale, CFI = .95) [Caplan(2010)]. We adapted six items from the scale to specifically reference chatbot use, maintaining representation of subscales measuring compulsive use, cognitive preoccupation, and negative outcomes. Existing work shows that loneliness is both a cause and effect of problematic internet use [Kim et al.(2009)], and the GPIUS2 was developed with online social interactions in mind, making it well-suited for examining potential compulsive or problematic patterns in human-AI interaction.
  • State Empathy Scale [Shen(2010)]: A validated measure capturing affective, cognitive, and associative dimensions of empathy. The scale demonstrates strong construct validity and reliability across multiple studies, with confirmatory factor analyses supporting its unidimensional structure [Shen(2010)]. Adjusted wording to be more general for application to both chatbots and people. Selected as a way to assess how participants interpersonally connect with chatbots, as research suggests people process parasocial relationships similarly to interpersonal ones [Schiappa et al.(2005)] and there may be parallels between parasocial relationships and human-chatbot relationships.
@@
  • Interpersonal Attraction Scale [McCroskey and McCain(1974)]: Measures three dimensions of attraction: social (desire for friendship), physical (appearance), and task (competence-based). The scale’s application to mediated relationships is supported by extensive parasocial interaction research [Schiappa et al.(2005)]. We shortened each subscale to four items based on factor loadings and created two versions: one for chatbots and one for close persons (excluding physical attraction for the latter due to potential family relationships).
@@
  • Perceived Homophily Scale [Mccroskey et al.(1975)]: Measures perceived similarity in communication, which research shows is central to both interpersonal and parasocial relationship development [Schiappa et al.(2005)]. Modified to use unidirectional Likert statements (e.g., "like me" instead of "unlike me – like me"). Selected 6 items based on factor loadings and created parallel versions for chatbots and a close person.
@@
  • Attributional Confidence Scale [Gudykunst and Nishida(1986)]: Assesses ability to predict others’ attitudes, feelings, and behavior [Clatterbuck(1979)]. Its relevance to human-AI interaction is supported by uncertainty reduction theory’s application to parasocial relationships; that is, relationships develop as people increase their ability to predict others’ behavior [Perse and Rubin(1989)]. Shortened to 5 items, removing items irrelevant to chatbots.
  • System Usability Scale (SUS) [Brooke(1996)]: A widely-validated measure of perceived system usability demonstrating high reliability ($\alpha$ = .91) across diverse interfaces [Bangor et al.(2008)]. Adjusted phrasing to refer specifically to chatbots. Chosen for consideration of potential links between usability and perceived closeness.
We acknowledge that modifying established scales may affect their validated psychometric properties. To manage survey length while maintaining measurement integrity, we shortened scales based on factor loadings from previous studies when available. While this approach allowed us to explore multiple potential factors, we recognize this as a limitation and address it through statistical controls for multicollinearity in our analysis, Benjamini-Hochberg correction for multiple comparisons, and conservative interpretation of results from modified scales. This exploratory approach allows us to identify potential factors warranting further investigation while maintaining statistical rigor through appropriate controls.